# Protect Before Generate: Error Correcting Codes within Discrete Deep Generative Models

### v1

@@ Abstract
Despite significant advancements in deep probabilistic models, learning low-dimensional discrete latent representations remains a challenging task. In this paper, we introduce a novel method that enhances variational inference in discrete latent variable models by leveraging Error-Correcting Codes to introduce redundancy in the latent representations. This redundancy is then exploited by the variational posterior to yield more accurate estimates, thereby narrowing the variational gap. Inspired by ECCs commonly used in digital communications and data storage, we demonstrate proof-of-concept using a Discrete Variational Autoencoder (DVAE) with binary latent variables and block repetition codes. We further extend this idea to a hierarchical structure based on polar codes, where certain latent bits are more robustly protected. Our method improves generation quality, data reconstruction, and uncertainty calibration compared to the uncoded DVAE, even when trained with tighter bounds such as the Importance Weighted Autoencoder (IWAE) objective. In particular, we demonstrate superior performance on MNIST, FMNIST, CIFAR10, and Tiny ImageNet datasets. The general approach of integrating ECCs into variational inference is compatible with existing techniques to boost variational inference, such as importance sampling or Hamiltonian Monte Carlo. We also outline the key properties ECCs must have to effectively enhance discrete variational inference.
Keywords: deep generative models, variational inference, discrete representations, error correcting codes
@@ 1 Introduction
Discrete latent space models seek to represent data using a finite set of features. Recent progress in generative models has increasingly favored these representations, as they are well-suited for datasets characterized by naturally discrete hidden states. However, effective learning of low-dimensional discrete latent representations is technically challenging. Vector Quantized-Variational Autoencoders [31, 21] stand out as solutions for this problem but rely on a non-probabilistic autoencoder, which does not provide uncertainty quantification in the latent space. To fit a fully probabilistic Variational Autoencoder (VAE) model [12], a common approach considers either Concrete [18] or Gumble-Softmax [11] approximations to sample from a discrete latent distribution in a reparameterizable manner [20, 17]. However, this approach leads to instabilities since the gradient variance is sensitive to the temperature that controls these approximations. The DVAE in [22, 30, 29], augments the binary latent representations with a set of continuous random variables, pairing each bit with a continuous counterpart where reparameterization can be done in a more stable manner after marginalizing the latent bits. The key distinction between the DVAE [22] and the DVAE++ [30] lies in their smoothing transformations: while [22] introduces spike-and-exponential transformations, [30] uses overlapping exponential distributions. These overlapping transformations are generalized in [29], enabling tighter variational bounds. We demonstrate the effectiveness of our method over a simplified version of the DVAE++ [30].
This work presents a novel method to improve variational inference and representation learning in generative models with discrete latent variables. In particular, for a latent variable model, we argue that one should use ECCs to introduce redundancy into the latent sample before the reconstruction decoder network processes it to generate the data. The variational approximation to the true posterior distribution can then exploit the added redundancy to provide more accurate estimates, reducing, in turn, the variational gap to the data likelihood.
Our approach is based on well-known digital communications and data storage techniques where information is protected with ECCs before transmission/storage to reduce the overall error rate during recovery. For different datasets, our results demonstrate that, compared to the uncoded DVAE, the DVAE with ECCs (Coded-DVAE) achieves superior generation quality, better data reconstruction, and critically calibrated uncertainty in the latent space. In Fig. 1, we highlight some representative results for both MNIST and FMNIST data sets. We note that the use of ECCs is a general design approach that is perfectly compatible with state-of-the-art techniques for improved variational inference, such as importance sampling [3, 27] or Hamiltonian Monte Carlo [32, 4]. In summary, our main contributions are:
  • We provide proof-of-concept results demonstrating that training deep generative models can be improved by ECC techniques, an idea that, to the best of our knowledge, is completely novel in the literature.
  • We formulate a coded version of DVAE using block repetition codes. We show that encoding/decoding of the block repetition code can be efficiently done with linear complexity.
  • We show that Coded-DVAE improves reconstruction, generation, and uncertainty calibration in the latent space when compared to the uncoded case using the same latent dimension, even when the uncoded DVAE is trained with tighter bounds such as the IWAE objective [3].
  • We discuss the generalization of this method to other coding schemes and introduce a hierarchical structure, inspired by polar codes [2], that effectively separates high-level information from finer details.
  • Through an extensive ablation study, we show that the enhancement in performance is not attributed to the increased dimensionality introduced by the redundancy from the ECC.
Figure 1: Comparison between uncoded and coded DVAE models with 8 latent bits, where the uncoded models are identified by the number of latent bits, and the coded models by their code rate. Fig. (a) presents examples of uncurated generation for FMNIST. Fig. (b) illustrates the evolution of the reconstruction and regularization terms of the loss. Fig. (c) depicts the graphical model of the uncoded DVAE, while Fig. (d) shows the graphical model of the coded DVAE.
@@ 2 Our baseline: the uncoded DVAE
This section introduces a simplified version of the uncoded DVAE [22, 30, 29], serving as the foundational model upon which the subsequent aspects of our work are constructed. Let $\boldsymbol{X}=\{\boldsymbol{x}_{0},\ldots,\boldsymbol{x}_{N}\}$ denote a collection of unlabelled data, where $\boldsymbol{x}_{i}$ represents a K-dimensional feature vector. While [22], [30] and [29] use Boltzmann machine priors, we consider a generative probabilistic model characterized by a simple low-dimensional binary latent variable $\boldsymbol{m}\in\{0,1\}^{M}$ comprising independent and identically distributed (i.i.d.) Bernoulli components $p(\boldsymbol{m})=\prod_{j=1}^{M}p(m_{j})=\prod_{j=1}^{M}\text{Ber}(\nu)$. Since backpropagation through discrete variables is generally not possible, a smoothing transformation of these binary variables is introduced. While the smoothing transformations proposed in [22] are limited to spike-and-X type of transformations, [30] show better results by using truncated exponential distributions:
$p(\boldsymbol{z}|\boldsymbol{m})=\prod_{j=1}^{M}p(z_{j}|m_{j}),\quad p(z_{j}|m_{j})$ $=\frac{e^{-\beta(z_{j}-m_{j})}}{Z_{\beta}},$ (1)
for $m_{j}\in\{0,1\}$, $z_{j}\in[0,1]$, and $Z_{\beta}=(1-e^{-\beta})/\beta$. The parameter $\beta$ serves as an inverse temperature term, similar to the one in the Gumbel-Softmax relaxation [11]. Given the simplicity of the defined binary prior, the complexity of the model is primarily determined by the likelihood function $p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})=p(f_{\boldsymbol{\theta}}(\boldsymbol{z}))$, where the likelihood is a Neural Network (NN) (referred to as the decoder) with parameter set $\boldsymbol{\theta}$.
Variational family and inference
Following [22], we assume an amortized variational family of the following form:
$q_{\boldsymbol{\eta}}(\boldsymbol{m},\boldsymbol{z}|\boldsymbol{x})$ $=q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})p(\boldsymbol{z}|\boldsymbol{m}),\quad q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})=\prod_{j=1}^{M}\text{Ber}(g_{j,\boldsymbol{\eta}}(\boldsymbol{x})),$ (2)
where $g_{\boldsymbol{\eta}}(\boldsymbol{x})$ represents a parameterized function; here, a NN (referred to as the encoder) with parameter set $\boldsymbol{\eta}$. Inference is achieved by maximizing the Evidence Lower Bound (ELBO), which can be expressed as
@@
where the first term corresponds to the reconstruction of the observed data and the second term is the Kullback-Leibler (KL) Divergence between the variational family and the binary prior distribution, which acts as a regularization term. This can be computed in closed form as $\mathcal{D}_{KL}\big{(}q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})||p(\boldsymbol{m})\big{)}=\sum_{j=1}^{M}\Big{[}q_{j}\log\frac{q_{j}}{\nu}+(1-q_{j})\log\frac{1-q_{j}}{1-\nu}\Big{]}$, where $q_{j}=q_{\boldsymbol{\eta}}(m_{j}=1|\boldsymbol{x})$. The reconstruction term needs to be approximated via Monte Carlo. Since $p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})$ does not depend on the binary latent variable $\boldsymbol{m}$, we can marginalize the posterior distribution as
$q_{\boldsymbol{\eta}}(\boldsymbol{z}|\boldsymbol{x})$ $=\prod_{j=1}^{M}q_{\boldsymbol{\eta}}(z_{j}|\boldsymbol{x}),\quad q_{\eta}(z_{j}|\boldsymbol{x})=\sum_{k=0}^{1}q_{\eta}(m_{j}=k|\boldsymbol{x})p(z_{j}|m_{j}=k).$ (4)
As shown in [30], the corresponding inverse Cumulative Density Function (CDF) is given by
@@
where $b=\big{(}\rho+e^{-\beta}(q_{j}-\rho)\big{)}/(1-q_{j})-1$ and $c=-[q_{j}e^{-\beta}]/(1-q_{j})$. The (5) is a differentiable function that converts a sample $\rho$ from an independent uniform distribution $\mathcal{U}(0,1)$ into a sample from $q_{\eta}(\boldsymbol{z}|\boldsymbol{x})$. Thus, we can apply the reparameterization trick to sample from the latent variable $\boldsymbol{z}$ and optimize the ELBO with respect to the model’s parameters. Algorithm 1 outlines the training procedure for the uncoded DVAE.
Algorithm 1 Training the model with uncoded inference.
1:  Input: training data $\boldsymbol{x}_{i}$.
2:  repeat
4:     $\boldsymbol{z}\leftarrow$ sample from (5)
6:     Compute ELBO according to (3)
7:     $\boldsymbol{\theta},\boldsymbol{\eta}\leftarrow Update(ELBO)$
8:  until convergence
3 Improving inference by adding redundancy to latent vectors
In the DVAE framework [22, 30, 29], the authors use Boltzmann machines as priors instead of the independent prior $p(\boldsymbol{m})$ presented in Section 2. While these complex priors increase the model’s flexibility and can produce competitive results, our objective is to enhance inference through model design by maintaining the simpler independent prior. This approach would improve interpretability and encourage the model to learn independent components in the latent space, which is essential for capturing potentially disentangled representations.
VAEs [12] are often viewed as lossy compression models, where the goal is to minimize reconstruction error while imposing regularization through a prior distribution. However, our approach is better understood from a generative standpoint. We first sample a latent vector $\boldsymbol{m}$, generate an observation $\boldsymbol{x}$, and focus on minimizing the error rate when recovering $\boldsymbol{m}$ from $\boldsymbol{x}$. Achieving this requires the variational approximation to be sufficiently accurate. In fields where reliable data transmission or storage is important, introducing ECCs is a well-established approach to reduce the error rate when estimating a discrete source $\boldsymbol{m}$ transmitted through a noisy channel with output $\boldsymbol{x}$. Estimating $\boldsymbol{m}$ from $\boldsymbol{x}$ implies approximating the true and unknown posterior distribution $p(\boldsymbol{m}|\boldsymbol{x})$ with a proposed $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$. The gap between $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ and $p(\boldsymbol{m}|\boldsymbol{x})$ is precisely the variational gap. We propose employing ECCs to safeguard $\boldsymbol{m}$ with controlled and known redundancy that can be leveraged by the variational posterior $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ by design. This way, it is possible to reduce the mistakes committed when comparing $\boldsymbol{m}$ with samples drawn from $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$, obtaining a tighter approximation to the true posterior $p(\boldsymbol{m}|\boldsymbol{x})$, therefore reducing the gap to optimal inference.
ECCs play a crucial role in information theory and digital communications by enabling reliable data transmission over unreliable channels [19]. They introduce redundancy into the transmitted data, allowing the receiver to detect errors and, in many cases, correct them without retransmission. In his seminal work, Shannon [25] demonstrated that arbitrarily reliable communication is possible through error correction. Our approach builds on the idea that the generative model in Fig. 1(c) can be conceptualized as a communication system, where the bits sampled from $p(\boldsymbol{m})$ undergo continuous modulation into $\boldsymbol{z}$ and are then transmitted through a nonlinear communication channel (in this setting, the decoder NN) characterized by the input/output response $p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})=p(f_{\boldsymbol{\theta}}(\boldsymbol{z}))$. In this scenario, the complexity of the channel is essential since it is necessary to account for the intricate nature of the data at its output (e.g., complex images). Following this idea, the process of inference via $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ can be thought of as deciphering the latent variable $\boldsymbol{m}$ given the observed data $\boldsymbol{x}$, where the encoder NN plays the role of the channel equalizer, trying to reverse the channel’s effects without knowing the bit correlations from the ECC.
4 Coded DVAE

### v2

# Improved Variational Inference in Discrete VAEs using Error Correcting Codes

@@ Abstract
Despite advances in deep probabilistic models, learning discrete latent representations remains challenging. This work introduces a novel method to improve inference in discrete Variational Autoencoders by reframing the inference problem through a generative perspective. We conceptualize the model as a communication system, and propose to leverage Error Correcting Codes (ECCs) to introduce redundancy in latent representations, allowing the variational posterior to produce more accurate estimates and reduce the variational gap. We present a proof-of-concept using a Discrete Variational Autoencoder with binary latent variables and low-complexity repetition codes, extending it to a hierarchical structure for disentangling global and local data features. Our approach significantly improves generation quality, data reconstruction, and uncertainty calibration, outperforming the uncoded models even when trained with tighter bounds such as the Importance Weighted Autoencoder objective. We also outline the properties that ECCs should possess to be effectively utilized for improved discrete variational inference.
@@ 1 Introduction
Discrete latent space models represent data using a finite set of features, making them particularly well-suited for representing inherently discrete data modalities. Additionally, they provide benefits in terms of interpretability and computational efficiency compared to continuous representations [Van Den Oord et al., 2017, Jang et al., 2017, Vahdat et al., 2018a]. Consequently, recent advancements in deep generative models have increasingly embraced discrete latent representations [Razavi et al., 2019, Ramesh et al., 2021, Rombach et al., 2022].
However, model training and inference of low-dimensional discrete latent representations remains technically challenging due to non-differentiability, which often requires the use of approximations that can lead to unstable optimization [Jang et al., 2017, Kool et al., 2019]. Currently, Vector Quantized-Variational Autoencoders [Van Den Oord et al., 2017, Razavi et al., 2019] stand out as state-of-the-art discrete latent variable models. They rely on a non-probabilistic encoder trained with a straight-through estimator, which lacks uncertainty quantification in the latent space. In contrast, fully probabilistic discrete Variational Autoencoders [Kingma and Welling, 2014] employ continuous relaxations such as Gumbel-Softmax [Jang et al., 2017] or Concrete [Maddison et al., 2017] for differentiable sampling [Ramesh et al., 2021, Lievin et al., 2020]. However, these methods can be unstable, as gradient variance is highly sensitive to the temperature parameter. The Discrete Variational Autoencoder (DVAE) framework [Rolfe, 2016, Vahdat et al., 2018a, b] addresses this issue by augmenting binary latent representations with continuous variables, allowing for stable reparameterization after marginalizing the latent bits. In this setting, Boltzmann machines serve as discrete priors to enhance model flexibility, at the cost of degraded interpretability.
This work introduces a novel approach to improve variational inference in discrete latent variable models, particularly discrete VAEs, by offering an alternative perspective on the inference problem. Rather than viewing VAEs as lossy compression methods employed to minimize data reconstruction error, we adopt a generative perspective, where inference seeks to recover the latent vector that generated a given observation. This allows us to conceptualize the generative model as a communication system, where we reinterpret latent vectors as messages transmitted through a nonlinear channel given by the VAE decoder, and then recovered by the VAE encoder. Based on this analogy, we propose to incorporate Error Correcting Codes to introduce redundancy into latent representations before they are processed by the decoder to generate data.
Figure 1: ECC within a DVAE. Fig. (a) illustrates the method and how adding redundancy transforms a dense $M$-dimensional latent space into a sparse $D$-dimensional one with $2^{M}$ valid vectors, enabling error correction in the variational posterior via minimum distance. Fig. (b) compares the ELBO and IWAE objectives for coded and uncoded models, where coding schemes with $M=8$ and $D=80,160,240$ introduce $D/8$ redundancy bits per latent bit.
Our approach builds on well-established techniques from digital communications and data storage, where information is protected using ECCs before transmission or storage to reduce the overall error rate during recovery [Moon, 2005]. Notably, Shannon’s landmark work demonstrated that properly designed error correction schemes can achieve arbitrarily low error rates [Shannon, 1948]. We use ECCs to deterministically introduce controlled redundancy to the latent vectors, increasing their Hamming distance and creating a higher-dimensional but sparser space where only a subset of vectors are valid. As illustrated in Fig. 1(a), this increased separation between latent codes allows for some errors in inference while still enabling recovery of the correct latent code by identifying the closest valid vector. Hence, this structure facilitates error correction during inference.
The integration of ECCs into discrete latent variable models is a novel approach that remains compatible with existing training methods for discrete generative models. Note that, while deep learning has been widely explored in digital communications [Guo et al., 2022, Wu et al., 2023, Shen et al., 2023, Ye et al., 2024, Chen et al., 2024], the use of ECCs as a design tool in machine learning remains largely unexplored. This work introduces a new research direction, leveraging communications theory to enhance deep generative models.
In our experiments, we use a simplified version of DVAE++ [Vahdat et al., 2018b] (referred to as uncoded DVAE) and show that the added redundancy enables a more accurate variational approximation to the true posterior, using simple independent priors. We highlight that error rates and variational gaps are linked through bounds derived from mismatch hypothesis testing, showing that minimizing the variational gap tightens the upper bound on the error rate [Schlüter et al., 2013, Yang et al., 2024]. Across different datasets, our results indicate that the DVAE with ECCs (Coded-DVAE) leads to reduced error rates in inference, resulting in smaller variational gaps. This improvement translates into superior generation quality, improved data reconstruction, and critically calibrated uncertainty in the latent space. Notably, Fig. 1(b) shows that Coded-DVAE achieves a significantly tighter training bounds even when the baseline is trained using the Importance Weighted Autoencoder (IWAE) objective [Burda et al., 2016].
Contributions. The primary contribution of this work is presenting a new perspective on the inference problem by framing the generative model as a communication system. We present proof-of-concept results that show how training discrete latent variable models can be improved by incorporating ECC techniques, an approach that, to the best of our knowledge, is novel in the existing literature. We then introduce a coded version of a discrete VAE, utilizing block repetition codes, and show that encoding and decoding can be done with linear complexity. Our results show that the Coded-DVAE model significantly reduces error rates during inference, leading to better data reconstruction, generation quality, and improved uncertainty calibration in the latent space. We further generalize this approach to other coding schemes and propose a hierarchical structure inspired by polar codes [Arikan, 2009], which effectively separates high-level information from finer details.  [^1] Code: https://github.com/mariamartinezgarcia/codedVAE
@@ 2 Our baseline: the uncoded DVAE
This section introduces a simplified version of the DVAE++ [Vahdat et al., 2018b] (uncoded DVAE), which serves as the foundational model upon which the subsequent aspects of our work are constructed. A key advantage of this framework is its stable training process, facilitated by reparameterization through smoothing transformations, in contrast to more unstable methods like Gumbel-Softmax [Jang et al., 2017]. To ensure that any performance improvements in the Coded-VAE stem solely from the coding scheme rather than a complex prior, we adopt a simple independent prior.
Figure 2: Graphical model of the uncoded DVAE.
Let $\boldsymbol{X}=\{\boldsymbol{x}_{0},\ldots,\boldsymbol{x}_{N}\}$ denote a collection of unlabeled data, where $\boldsymbol{x}_{i}\in\mathbb{R}^{K}$ represents a $K$-dimensional feature vector. While Rolfe [2016], Vahdat et al. [2018b] and Vahdat et al. [2018a] use Boltzmann machine priors, we consider a generative probabilistic model characterized by a simple low-dimensional binary latent variable $\boldsymbol{m}\in\{0,1\}^{M}$ comprising independent and identically distributed (i.i.d.) Bernoulli components $p(\boldsymbol{m})=\prod_{j=1}^{M}p(m_{j})=\prod_{j=1}^{M}\text{Ber}(\nu)$, where $\nu=p(m_{j}=1)$. Since backpropagation through discrete samples is generally not possible, a smoothing transformation $p(\boldsymbol{z}|\boldsymbol{m})=\prod_{j=1}^{M}p(z_{j}|m_{j})$ of these binary variables is introduced. Following Vahdat et al. [2018b], we introduce truncated exponential distributions given by
$p(z|m=1)=\frac{e^{\beta(z-1)}}{Z_{\beta}},\quad p(z|m=0)=\frac{e^{-\beta z}}{Z_{\beta}},$ (1)
for $m\in\{0,1\}$, $z\in[0,1]$, and $Z_{\beta}=(1-e^{-\beta})/\beta$. The parameter $\beta$ serves as an inverse temperature term, similar to the one in the Gumbel-Softmax relaxation [Jang et al., 2017]. Given the simplicity of the defined binary prior, the complexity of the model is primarily determined by the likelihood function $p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})=p(f_{\boldsymbol{\theta}}(\boldsymbol{z}))$, the decoder function, where the likelihood is a Neural Network (NN) (referred to as the decoder) with parameter set $\boldsymbol{\theta}$.
2.1 Variational family and inference
Following Rolfe [2016], we assume an amortized variational family of the form $q_{\boldsymbol{\eta}}(\boldsymbol{m},\boldsymbol{z}|\boldsymbol{x})=q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})p(\boldsymbol{z}|\boldsymbol{m})$, where
$q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ $=\prod_{j=1}^{M}\text{Ber}(m_{j};q_{j})=\prod_{j=1}^{M}\text{Ber}(m_{j};g_{j,\boldsymbol{\eta}}(\boldsymbol{x}))$ (2)
where $\text{Ber}(m_{j};q_{j})$ is a Bernoulli with parameter $q_{j}=q_{\boldsymbol{\eta}}(m_{j}=1|\boldsymbol{x})$ and $g_{\boldsymbol{\eta}}(\boldsymbol{x})$ is a NN (referred to as the data encoder) with parameter set $\boldsymbol{\eta}$. Inference is achieved by maximizing the Evidence Lower Bound (ELBO), which can be expressed as
@@ 2.1 Variational family and inference
where the first term corresponds to the reconstruction of the observed data and the second is the Kullback-Leibler (KL) Divergence between the variational family and the prior distribution, which acts as a regularization term. This can be computed in closed form as $\mathcal{D}_{KL}\big{(}q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})||p(\boldsymbol{m})\big{)}=\sum_{j=1}^{M}\Big{[}q_{j}\log\frac{q_{j}}{\nu}+(1-q_{j})\log\frac{1-q_{j}}{1-\nu}\Big{]}$, with $\nu=p(m_{j}=1)$. Since $p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})$ does not depend on the binary latent variable $\boldsymbol{m}$, we can marginalize the posterior distribution as
$\begin{gathered}q_{\boldsymbol{\eta}}(\boldsymbol{z}|\boldsymbol{x})=\prod_{j=1}^{M}q_{\boldsymbol{\eta}}(z_{j}|\boldsymbol{x}),\\ \quad q_{\eta}(z_{j}|\boldsymbol{x})=\sum_{k=0}^{1}q_{\eta}(m_{j}=k|\boldsymbol{x})p(z_{j}|m_{j}=k).\end{gathered}$ (4)
As shown in Vahdat et al. [2018b], the corresponding inverse Cumulative Density Function (CDF) is given by
@@
where $b=\big{(}\rho+e^{-\beta}(q_{j}-\rho)\big{)}/(1-q_{j})-1$ and $c=-[q_{j}e^{-\beta}]/(1-q_{j})$. Equation (5) defines a differentiable function that converts a sample $\rho$ from an independent uniform distribution $\mathcal{U}(0,1)$ into a sample from $q_{\boldsymbol{\eta}}(\boldsymbol{z}|\boldsymbol{x})$. Thus, we can apply the reparameterization trick to sample from the latent variable $\boldsymbol{z}$ and optimize the ELBO.
3 On the variational gap and the inference error
While VAEs are typically viewed as lossy compression methods, this work introduces an alternative perspective on the inference problem from a generative viewpoint: we first sample a latent vector $\boldsymbol{m}$, generate an observation $\boldsymbol{x}$, and aim to minimize the error rate when recovering $\boldsymbol{m}$ from $\boldsymbol{x}$. This requires approximating the true, unknown posterior distribution $p_{\boldsymbol{\theta}}(\boldsymbol{m}|\boldsymbol{x})$ with a sufficiently accurate proposed distribution $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$. The gap between $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ and $p_{\boldsymbol{\theta}}(\boldsymbol{m}|\boldsymbol{x})$ is precisely the variational gap, which can be related to the error rate in inference, i.e. comparing the true $\boldsymbol{m}$ with samples drawn from $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$.
In statistical classification, particularly in the context of multiple hypothesis testing, classification error is the primary performance metric. Bayes’ decision rule (also known as Bayes’ test) minimizes classification error, assuming the true probability distributions governing the classification problem are known. In the case of model mismatch, e.g., when a posterior distribution is learned using variational inference, statistical bounds can be used to relate the classification error and the model estimation error [Schlüter et al., 2013, Yang et al., 2024]. In particular, for a generative model $\boldsymbol{x}\sim p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{m})$ with discrete latent vector $\boldsymbol{m}$, let $1-A_{\boldsymbol{\theta}}$ be the error rate when estimating $\boldsymbol{m}$ from $\boldsymbol{x}$ using the true posterior $p_{\boldsymbol{\theta}}(\boldsymbol{m}|\boldsymbol{x})$ and $1-A_{\boldsymbol{\eta}}$ be the error rate using the variational approximate posterior $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$. Then, $\Delta\doteq A_{\boldsymbol{\theta}}-A_{\boldsymbol{\eta}}$ can be bounded as
$0\leq\Delta^{2}$ $\leq 1-e^{-2\mathcal{D}_{KL}(q_{\boldsymbol{\eta}}(\boldsymbol{m},\boldsymbol{x})||p_{\boldsymbol{\theta}}(\boldsymbol{m},\boldsymbol{x}))},$ (6)
where $\mathcal{D}_{KL}(q_{\boldsymbol{\eta}}(\boldsymbol{m},\boldsymbol{x})||p_{\boldsymbol{\theta}}(\boldsymbol{m},\boldsymbol{x}))=\mathbb{E}_{p(\boldsymbol{x})}[\mathcal{D}_{KL}(q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})||p_{\boldsymbol{\theta}}(\boldsymbol{m}|\boldsymbol{x}))]$ is the variational gap. For fixed $\boldsymbol{\theta}$ (and thus for fixed $A_{\boldsymbol{\theta}}$), any improvement in $\boldsymbol{\eta}$ that reduces the expected variational gap also tightens an upper bound on $\Delta\doteq A_{\boldsymbol{\theta}}-A_{\boldsymbol{\eta}}$.
In fields where reliable data transmission or storage is crucial, introducing ECCs is a well-established approach to reduce the error rate when estimating a discrete source $\boldsymbol{m}$ transmitted through a noisy channel with output $\boldsymbol{x}$. Building on this, we propose using ECCs to protect $\boldsymbol{m}$ with controlled redundancy that the variational posterior $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ can leverage. This way, it is possible to reduce error rates in inference and refine the variational parameters $\boldsymbol{\eta}$ on the enhanced model, thus narrowing the variational gap. Notably, according to (6), a smaller variational gap results in a tighter upper bound on the error rates. In the following sections, we show that simple coding schemes enhance inference, improving generation, reconstruction, log-likelihoods, and uncertainty calibration in the latent space.
4 Coded-DVAE
Figure 3: Graphical model of the Coded-DVAE.
Model BER BER MAP WER WER MAP LL test
uncoded 5 0.051 0.051 0.195 0.190 -267.703
[…]
uncoded 8 0.089 0.088 0.384 0.376 -249.880
coded 8/80 0.021 0.003 0.144 0.021 -232.992
coded 8/160 0.027 0.001 0.189 0.009 -235.819
coded 8/240 0.037 0.004 0.231 0.023 -238.459
uncoded 10 0.142 0.144 0.622 0.644 -244.997
coded 10/100 0.040 0.005 0.321 0.036 -230.772
coded 10/200 0.044 0.006 0.341 0.039 -234.849
coded 10/300 0.045 0.002 0.349 0.014 -238.647
Figure 4: Evaluation of generation in FMNIST. Uncurated randomly generated samples (left) and evaluation metrics (right), including BER and WER sampled from the variational posterior, BER and WER using the MAP value, and test LL.